OpenStudy (anonymous):

What is the indefinite integral of (1 - x^2)/x

OpenStudy (anonymous):

Simplify the integrand to: \(\large \int \frac{1-x^2}{x}dx=\int (\frac{1}{x}-\frac{x^2}{x}) dx=\int (\frac{1}{x}-x)dx \)

OpenStudy (anonymous):

Now that should be doable...

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

you're welcome :)